OpenStudy (anonymous):

please help! I can't figure out the two x's in this equation because I'm not sure I can simplify it beforehand. The equation is in the comments...

OpenStudy (anonymous):

\[\frac{ x }{ x-12 } = \frac{ 1 }{ x-1 }\]

OpenStudy (anonymous):

"cross multiply" as they say, and start with \[x(x-1)=x-12\]

OpenStudy (anonymous):

then multiply out on the left to get \[x^2-x=x-12\]

OpenStudy (anonymous):

put all on one side of the equal sign by subtracting \(x\) and adding \(12\) and get \[x^2-2x+12=0\] then try to factor or use the quadratic formula to solve for \(x\)

OpenStudy (anonymous):

viola! standard form.

OpenStudy (anonymous):

i think the problem is that the solutions are not real numbers

OpenStudy (anonymous):

awesome! I get it from that part (: thank you so much! and yes, it's a multiple choice question and the answers aren't real numbers. So it's fine!

OpenStudy (anonymous):

oh in that case you should not set it equal to zero leave it as \[x^2-2x=-12\] complete the square via \[(x-1)^2=-12+1=-11\] and get \[x=1\pm\sqrt{11}i\] more quickly without formula

OpenStudy (anonymous):

oh okay, thank you so much! I appreciate your help!

OpenStudy (anonymous):

yw